OpenStudy (anonymous):

The functions f(x) and g(x) are described below: f(x) = 2x + 9 g(x) = 2x - 15 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

would it be 15 @Nnesha

OpenStudy (aum):

In y = mx + b, 'b' is the y-intercept. What is the y-intercept of f(x) = 2x + 9 ? What is the y-intercept of g(x) = 2x - 15 ?

OpenStudy (anonymous):

9 and -15

OpenStudy (anonymous):

It looks like its asking the difference between the y-intercepts

OpenStudy (aum):

correct. So f(x) has a y-intercept of 9. How much down does f(x) have to be shifted to have a y-intercept of -15?

OpenStudy (anonymous):

has to be shifted down 15

OpenStudy (aum):

f(x) has a y-intercept of 9. If you shift it down by 9 units, it will have a y-intercept of 0. After that you have to further shift it down by 15 units to have a y-intercept of -15. so 9 + 15 = 24 units.

OpenStudy (aum):

Or simply, 9 - (-15) = 9 + 15 = 24 units.

OpenStudy (anonymous):

oooh sorry my bad

OpenStudy (anonymous):

thanks!

OpenStudy (aum):

yw.